### v2

# Higher Jacobian matrix of weighted homogeneous polynomials and derivation algebras

Abstract.
We prove that the ideal generated by the maximal minors of the higher-order Jacobian matrix of a weighted homogeneous polynomial is also weighted homogeneous. As an application, we give a partial answer to a conjecture concerning the non-existence of negative weight derivations on the higher Nash blowup local algebra of a hypersurface.
Key words and phrases:
Higher Nash blowup local algebras, higher-order Jacobian matrix, weighted homogeneous isolated hypersurface singularities
2020 Mathematics Subject Classification:
14B05, 32S05, 13N15.
Introduction
The Jacobian matrix of order $n$ is a higher-order version of the classical Jacobian matrix. It was introduced as a tool for computing the higher Nash blowup of a hypersurface [Dua17]. Higher-order Jacobian matrices were later rediscovered and further developed by several authors [BJNB19, BD20, CMDGF21]. Ever since, this matrix has seen a wide variety of applications: in the study of higher Nash blowups [Dua17, CMDGF21], the study of invariants of rings in positive characteristic [BJNB19], the study of homological properties of the module of higher Kähler differentials [BD20, dAD21], relations with Hasse-Schmidt derivations and jet schemes [LDS24, Bar23], invariants of germs of hypersurface singularities [HMYZ23, LY25], motivic zeta functions [LY25]. In addition to these applications, new questions has been raised regarding general properties of higher Jacobian matrices.
Let $f\in\mathbb{C}\{x_{1},\ldots,x_{s}\}$, where $\mathbb{C}\{x_{1},\ldots,x_{s}\}$ denotes the ring of convergent power series at the origin. Let $\operatorname{Jac}_{n}(f)$ be the higher Jacobian matrix of $f$ and $\mathcal{J}_{n}(f)$ the ideal generated by the maximal minors of $\operatorname{Jac}_{n}(f)$ (see Definition 1.3). N. Hussain, G. Ma, S. S.-T. Yau, and H. Zuo defined the higher Nash blowup local algebra as the quotient $\mathcal{T}_{n}(f)=\mathbb{C}\{x_{1},\ldots,x_{s}\}/\langle f,\mathcal{J}_{n}(f)\rangle$ [HMYZ23]. For $n=1$, this is the classical Tjurina algebra. The authors proposed several conjectures regarding $\mathcal{T}_{n}(f)$: invariance under contact equivalence, homogeneity properties, and bounds for their dimensions. They also provided some evidence that supports the conjectures. These questions can be seen as higher-order analogues of classical results in singularity theory. For instance, a famous result by J. Mather and S. S.-T. Yau states that the Tjurina algebra is a complete invariant of isolated hypersurface singularities under contact equivalence [MY82]. Moreover, higher-order versions of the Tjurina algebra have been an object of intense study (see, for instance, [GLS07, DS15, GP17, DGI20, MYZ23]).
Very recently, the conjecture regarding the invariance of $\mathcal{T}_{n}(f)$ under contact equivalence was proved by Q. T. Lê and T. Yasuda [LY25]. In this paper we are interested in exploring the questions regarding the homogeneity properties of the higher Jacobian matrix. The following is our first main theorem.
Theorem A (see Theorem 1.7).
Let $f\in\mathbb{C}[x_{1},\ldots,x_{s}]$ be a weighted homogeneous polynomial. Then $\mathcal{J}_{n}(f)$ is a weighted homogeneous ideal for every $n\in\mathbb{N}$.
Another object introduced by the aforementioned authors is the algebra of derivations $\mathcal{L}_{n}(f)=\operatorname{Der}(\mathcal{T}_{n}(f))$. For $n=1$, this is known as the Yau algebra of $f$. A classical result in singularity theory states that the Yau algebra is a solvable Lie algebra for isolated hypersurface singularities [Yau86, Yau91]. Moreover, the Yau algebra has been used to distinguish complex analytic structures of isolated hypersurface singularities [SY90].
By Theorem A, $\mathbb{C}[x_{1},\ldots,x_{s}]/\langle f,\mathcal{J}_{n}(f)\rangle$ is a graded algebra whenever $f$ is a weighted homogeneous polynomial. In this case, the algebra $\operatorname{Der}(\mathbb{C}[x_{1},\ldots,x_{s}]/\langle f,\mathcal{J}_{n}(f)\rangle)$ is also graded. There is a general conjecture on algebras of derivations of local Artinian graded algebras, known as Halperin conjecture, stating the non-existence of negative weight derivations over those algebras [Che99]. This conjecture has been intensively studied [Mei82, FHT01, CYZ19, CCYZ20, CHYZ20]. Inspired by Halperin’s question, N. Hussain, G. Ma, S. S.-T. Yau, and H. Zuo conjectured that there are no negative weight derivations on $\mathbb{C}[x_{1},\ldots,x_{s}]/\langle f,\mathcal{J}_{n}(f)\rangle$. They proved their conjecture in the case $n=s=2$. Our second main result shows that the conjecture is true for $n\geq 3$ and $s=2$.
Theorem B (see Theorem 2.6).
Let $f\in\mathbb{C}[x_{1},x_{2}]$ be a weighted homogeneous polynomial of weight $(w_{1},w_{2})\in\mathbb{N}^{2}$ and degree $d$ that defines an isolated hypersurface singularity. Suppose that $d\geq 2w_{1}\geq 2w_{2}>0$. Let $n\geq 3$. Then $\operatorname{Der}(\mathbb{C}[x_{1},x_{2}]/\langle f,\mathcal{J}_{n}(f)\rangle)$ is non-negatively graded.
This paper is divided as follows. In the first section we prove Theorem A. We also establish some facts regarding the weighted degree of the maximal minors of the higher Jacobian matrix. The second section contains the proof of Theorem B.
1. Higher-order Jacobian matrix of a weighted homogeneous polynomial
The following notation will be constantly used throughout this paper.
Notation 1.1.
Let $n,s\in\mathbb{N}_{\geq 1}$. Given $\gamma=(\gamma_{1},\ldots,\gamma_{s})\in\mathbb{N}^{s}$, denote $|\gamma|=\gamma_{1}+\cdots+\gamma_{s}$. For $\gamma,\gamma^{\prime}\in\mathbb{N}^{s}$ we denote $\gamma\cdot\gamma^{\prime}$ the usual Euclidean inner product. In addition, consider
$B:=\{\beta\in\mathbb{N}^{s}|0\leq|\beta|\leq n-1\},$
$B_{1}:=\{\beta\in\mathbb{N}^{s}|1\leq|\beta|\leq n-1\},$
$A:=\{\alpha\in\mathbb{N}^{s}|0\leq|\alpha|\leq n\},$
$A_{1}:=\{\alpha\in\mathbb{N}^{s}|1\leq|\alpha|\leq n\},$
$C_{n}:=\{\alpha\in\mathbb{N}^{s}||\alpha|=n\}.$
Remark 1.2.
It is known that
$\operatorname{Card}(B)=\binom{s+n-1}{s}=:M,$
$\operatorname{Card}(A)=\binom{s+n}{s}=:N,$
$\operatorname{Card}(C_{n})=\binom{n+s-1}{s-1}=N-M=:l.$
Definition 1.3 ([Dua17, BJNB19, BD20]).
Let $f\in\mathbb{C}\{x_{1},\ldots,x_{s}\}$. Denote
$\operatorname{Jac}_{n}(f):=\left(\frac{1}{(\alpha-\beta)!}\frac{\partial^{\alpha-\beta}(f)}{\partial x^{\alpha-\beta}}\right)_{\begin{subarray}{c}\beta\in B\\ \alpha\in A_{1}\end{subarray}},$
where we define ${\frac{1}{(\alpha-\beta)!}\frac{\partial^{\alpha-\beta}(f)}{\partial x^{\alpha-\beta}}=0},$ whenever $\alpha_{i}<\beta_{i}$ for some $i$. It is a $M\times(N-1)$-matrix. We call $\operatorname{Jac}_{n}(f)$ the higher-order Jacobian matrix of $f$ or the Jacobian matrix of order $n$ of $f$. We order this matrix increasingly using a graded lexicographical order and taking $x_{1}<x_{2}<\cdots<x_{s}$. Moreover, denote as $\mathcal{J}_{n}(f)$ the ideal generated by all maximal minors of $\operatorname{Jac}_{n}(f)$.
Remark 1.4.
The higher-order Jacobian matrix was originally introduced as a tool to compute the higher Nash blowup of a hypersurface [Yas07, Dua17].
Example 1.5.
Let $f=x^{3}-y^{2}\in\mathbb{C}[x,y]$. Then
$\operatorname{Jac}_{2}(f)=\left(\begin{array}[]{ccccc}3x^{2}&-2y&3x&0&-1\\ f&0&3x^{2}&-2y&0\\ 0&f&0&3x^{2}&-2y\\ \end{array}\right).$
N. Hussain, G. Ma, S. S.-T. Yau, and H. Zuo proposed the following conjecture regarding homogeneity properties of $\operatorname{Jac}_{n}(f)$.
Conjecture 1.6 ([HMYZ23, Conjecture 1.7]).
Let $f\in\mathbb{C}[x_{1},\ldots,x_{s}]$ be a weighted homogeneous polynomial. Then $\mathcal{J}_{n}(f)$ is a weighted homogeneous ideal.
The aforementioned authors verified Conjecture 1.6 in the case $s=n=2$ [HMYZ23, Lemma 4.1]. The following theorem gives a positive answer to Conjecture 1.6 for arbitrary $n$ and $s$.
Theorem 1.7.
Let $f$ be a weighted homogeneous polynomial in $\mathbb{C}[x_{1},\ldots,x_{s}]$ of weight $w\in\mathbb{N}^{s}$ and degree $d$. Then $\mathcal{J}_{n}(f)$ is a weighted homogeneous ideal for every $n\in\mathbb{N}$.
Proof.
Let $L$ be a matrix formed by taking $M$ columns of $\operatorname{Jac}_{n}(f)$. Let $B=\{\beta(1),\ldots,\beta(M)\}$ and $\{\alpha(1),\ldots,\alpha(M)\}\subset A_{1}$ be the sets of vectors indexing the rows and columns of $L$, respectively. Denote the entries of $L$ as $L_{\beta(i)\alpha(j)}$ with $1\leq i,j\leq M$. Recall that
$L_{\beta(i)\alpha(j)}=\frac{1}{(\alpha(j)-\beta(i))!}\frac{\partial^{\alpha(j)-\beta(i)}(f)}{\partial x^{\alpha(j)-\beta(i)}}.$
[…]
$\det(L)=\sum_{\theta\in I}\operatorname{sgn}(\theta)L_{\beta(1)\alpha(\theta(1))}\cdots L_{\beta(M)\alpha(\theta(M))}.$
Notice that $f$ being weighted homogeneous of degree $d$ implies that $L_{\beta(i)\alpha(\theta(i))}$ is a weighted homogeneous polynomial of degree $d-(\alpha(\theta(i))-\beta(i))\cdot w$. Hence, $L_{\beta(1)\alpha(\theta(1))}\cdots L_{\beta(M)\alpha(\theta(M))}$ is a weighted homogeneous polynomial of degree $dM-\sum_{i=1}^{M}(\alpha(\theta(i))-\beta(i))\cdot w$.
For every $\sigma\in I$ we claim that
$\sum_{i=1}^{M}(\alpha(\sigma(i))-\beta(i))\cdot w=\sum_{i=1}^{M}(\alpha(i)-\beta(i))\cdot w.$
Indeed,
0 0 $=\sum_{i=1}^{M}(\alpha(\sigma(i))-\alpha(i))\cdot w$
$=\sum_{i=1}^{M}(\alpha(\sigma(i))-\beta(i)+\beta(i)-\alpha(i))\cdot w$
$=\sum_{i=1}^{M}(\alpha(\sigma(i))-\beta(i))\cdot w-\sum_{i=1}^{M}(\alpha(i)-\beta(i))\cdot w.$
Taking $c=\sum_{i=1}^{M}(\alpha(i)-\beta(i))\cdot w$, we conclude $L_{\beta(1)\alpha(\theta(1))}\cdots L_{\beta(M)\alpha(\theta(M))}$ is a weighted homogeneous polynomial of weight $w$ and degree $dM-c$ for every $\theta\in I$. Therefore, $\det(L)$ is a weighted homogeneous polynomial of weight $w$ and degree $dM-c$. ∎
Notice that the proof of the previous theorem also exhibited the degree of the maximal minors of $\operatorname{Jac}_{n}(f)$. Our next goal is to give lower bounds for those degrees, in some special cases.
1.1. Homogeneous case
Let $f$ be a homogeneous polynomial in $\mathbb{C}[x_{1},\ldots,x_{s}]$ of degree $d$, and let $n\geq 2$, $s\geq 2$. Let $g$ be the determinant of the submatrix of $\operatorname{Jac}_{n}(f)$ formed by the columns indexed by $\{\alpha(1),\ldots,\alpha(M)\}\subset A_{1}$ and the rows indexed by $B=\{\beta(1),\ldots,\beta(M)\}$. From the proof of Theorem 1.7, $g$ is a homogeneous polynomial of degree $dM-c$, where
$c$ $=\sum_{i=1}^{M}|\alpha(i)-\beta(i)|=\sum_{i=1}^{M}|\alpha(i)|-\sum_{i=1}^{M}|\beta(i)|.$
Lemma 1.8.
With the previous notation we have $sM\geq c+s$.
Proof.
We can assume that $1\leq|\alpha(1)|\leq\cdots\leq|\alpha(M)|\leq n$. Recall the notation from Remark 1.2 and denote $\mathcal{A}:=\{\alpha(1),\ldots,\alpha(M-l)\}$. Notice that $\mathcal{A}\subset B_{1}$, thus
$\sum_{i=1}^{M}|\alpha(i)|$ $\leq\sum_{\alpha\in\mathcal{A}}|\alpha|+\sum_{\alpha\in C_{n}}|\alpha|=\sum_{\alpha\in\mathcal{A}}|\alpha|+ln=\sum_{\alpha\in\mathcal{A}}|\alpha|+sM.$
On the other hand, $\operatorname{Card}(B_{1}\setminus\mathcal{A})=(M-1)-(M-l)=l-1\geq s$. This implies
$\sum_{i=1}^{M}|\beta(i)|=\sum_{\beta\in B_{1}}|\beta|\geq\sum_{\alpha\in\mathcal{A}}|\alpha|+s.$
From the above,
$c+s$ $=\sum_{i=1}^{M}|\alpha(i)|-\sum_{i=1}^{M}|\beta(i)|+s$
$\leq sM+\sum_{\alpha\in\mathcal{A}}|\alpha|-\sum_{i=1}^{M}|\beta(i)|+s$
$\leq sM.$
Corollary 1.9.
Let $f\in\mathbb{C}[x_{1},\ldots,x_{s}]$ be a homogeneous polynomial of degree $d\geq s$ and let $n\geq 2$. Then $d$ is a lower bound for the degrees of the maximal minors of $\operatorname{Jac}_{n}(f)$. In other words, $d\geq s$ implies $dM-c\geq d$.
Proof.
By Lemma 1.8, $d(M-1)-c\geq s(M-1)-c\geq 0$. ∎
1.2. Weighted homogeneous case
[…]
As in the previous subsection, we want to give a lower bound for the weighted degrees of the maximal minors of $\operatorname{Jac}_{n}(f)$. Because we are working with an arbitrary weight, the strategy for finding such bound is not as straightforward as the ones of Lemma 1.8 and Corollary 1.9.
Lemma 1.10.
Let $w=(w_{1},w_{2})\in\mathbb{N}^{2}$. Then $\sum_{\alpha\in C_{n}}\alpha\cdot w=|w|M$.
Proof.
We have that
$\sum_{\alpha\in C_{n}}\alpha\cdot w$ $=\sum_{i=0}^{n}(n-i,i)\cdot(w_{1},w_{2})$
$=\sum_{i=0}^{n}(n-i)w_{1}+\sum_{i=0}^{n}iw_{2}$
$=\frac{n(n+1)}{2}w_{1}+\frac{n(n+1)}{2}w_{2}$
$=|w|M.$
Before going any further, we present an example that illustrates the notation and the key idea we develop to find the desired bound.
Example 1.11.
Let $n=3$. Thus $M=6$. In this case,
[…]
Consider the submatrix of $\operatorname{Jac}_{3}(f)$ defined by the columns
$\mathfrak{A}=\{(1,0),(0,1),(2,0),(1,1),(3,0),(1,2)\}.$
Let us show how we bound the sums $\sum_{\alpha\in\mathfrak{A}}\alpha\cdot w$ and $\sum_{\beta\in B}\beta\cdot w$.
Let $C=\mathfrak{A}\setminus C_{3}=\{(1,0),(0,1),(2,0),(1,1)\}$. Let $\theta=(1,0)$ and $\tau=(2,1)$. Notice that $\theta\in C$, $\tau\in C_{3}\setminus\mathfrak{A}$, and $\theta\cdot w<\tau\cdot w$. This implies
$\sum_{\alpha\in\mathfrak{A}}\alpha\cdot w$ $=\sum_{\alpha\in C}\alpha\cdot w+\sum_{\alpha\in\mathfrak{A}\cap C_{3}}\alpha\cdot w$
$=\sum_{\alpha\in C\setminus\{\theta\}}\alpha\cdot w+\theta\cdot w+\sum_{\alpha\in\mathfrak{A}\cap C_{3}}\alpha\cdot w$
$<\sum_{\alpha\in C\setminus\{\theta\}}\alpha\cdot w+\tau\cdot w+\sum_{\alpha\in\mathfrak{A}\cap C_{3}}\alpha\cdot w$
$<\sum_{\alpha\in C\setminus\{\theta\}}\alpha\cdot w+\sum_{\alpha\in C_{3}}\alpha\cdot w$
$=\sum_{\alpha\in C\setminus\{\theta\}}\alpha\cdot w+|w|M.$
On the other hand, we also have
$\sum_{\beta\in B}\beta\cdot w=\sum_{\beta\in B_{1}}\beta\cdot w$ $=\Big{(}\sum_{\begin{subarray}{c}\beta\in B_{1}\\ \beta\neq(1,0),(0,2)\end{subarray}}\beta\cdot w\Big{)}+(1,0)\cdot w+(0,2)\cdot w$
$=\Big{(}\sum_{\alpha\in C\setminus\{\theta\}}\alpha\cdot w\Big{)}+(1,0)\cdot w+(0,2)\cdot w$
$>\sum_{\alpha\in C\setminus\{\theta\}}\alpha\cdot w+w_{1}+w_{2}.$
The following lemma generalizes the process of the previous example.
Lemma 1.12.
Let $w=(w_{1},w_{2})\in\mathbb{N}^{2}$. Assume $w_{1}\geq w_{2}>0$. Let $\mathfrak{A}=\{\alpha(1),\ldots,\alpha(M)\}\subset A_{1}$, $B=\{\beta(1),\ldots,\beta(M)\}$, and $C=\mathfrak{A}\setminus C_{n}$. Define subsets $\mathcal{A}\subset C$ satisfying the following conditions:
  1. (1)
    $\mathcal{A}=C$, $\operatorname{Card}(B_{1}\setminus\mathcal{A})\geq 2$ and there exists $\theta\in B_{1}\setminus\mathcal{A}$ such that $\theta_{1}\geq 1$.
  2. (2)
    $\mathcal{A}=C\setminus\{\theta\}$, where $\theta\in C$ is such that $\theta_{1}\geq 1$, $\operatorname{Card}(B_{1}\setminus\mathcal{A})\geq 2$, and there exists $\tau\in C_{n}\setminus\mathfrak{A}$ such that $\theta\cdot w\leq\tau\cdot w$.
  3. (3)
    $\mathcal{A}=C\setminus\{\theta,\theta^{\prime}\}$, where $\theta,\theta^{\prime}\in C$ are such that $\theta_{1}\geq 1$ and there exist $\tau,\tau^{\prime}\in C_{n}\setminus\mathfrak{A}$ such that $\theta\cdot w\leq\tau\cdot w$ and $\theta^{\prime}\cdot w\leq\tau^{\prime}\cdot w$.
Then $|w|M\geq c+|w|$, where $c=\sum_{i=1}^{M}(\alpha(i)-\beta(i))\cdot w$.
Proof.
In each of the cases we have:
$\sum_{i=1}^{M}\alpha(i)\cdot w$ $=\sum_{\alpha\in C}\alpha\cdot w+\sum_{\alpha\in\mathfrak{A}\cap C_{n}}\alpha\cdot w$
$=\sum_{\alpha\in\mathcal{A}}\alpha\cdot w+\sum_{\alpha\in C\setminus\mathcal{A}}\alpha\cdot w+\sum_{\alpha\in\mathfrak{A}\cap C_{n}}\alpha\cdot w$
$\leq\sum_{\alpha\in\mathcal{A}}\alpha\cdot w+\sum_{\alpha\in C_{n}}\alpha\cdot w$
$=\sum_{\alpha\in\mathcal{A}}\alpha\cdot w+|w|M,$
where the last equality follows by Lemma 1.10.
On the other hand, assuming that $\beta(1)=(0,0)$, since $\operatorname{Card}(B_{1}\setminus\mathcal{A})\geq 2$ we also have:
$\sum_{i=1}^{M}\beta(i)\cdot w=\sum_{i=2}^{M}\beta(i)\cdot w\geq\sum_{\alpha\in\mathcal{A}}\alpha\cdot w+(w_{1}+w_{2}).$
Finally,
$c$ $=\sum_{i=1}^{M}\alpha(i)\cdot w-\sum_{i=1}^{M}\beta(i)\cdot w$
$\leq|w|M+\sum_{\alpha\in\mathcal{A}}\alpha\cdot w-\sum_{i=1}^{M}\beta(i)\cdot w$
$\leq|w|M+\sum_{\alpha\in\mathcal{A}}\alpha\cdot w-\sum_{\alpha\in\mathcal{A}}\alpha\cdot w-(w_{1}+w_{2})$
$=|w|M-|w|.$
Remark 1.13.
Let $f$ be a weighted homogeneous polynomial in $\mathbb{C}[x_{1},x_{2}]$ of weight $w\in\mathbb{N}^{2}$ and degree $d$. Without loss of generality we can assume that $w_{1}\geq w_{2}>0$. Let $n\geq 3$ and $g$ be the determinant of the submatrix of $\operatorname{Jac}_{n}(f)$ formed by the columns indexed by $\mathfrak{A}=\{\alpha(1),\ldots,\alpha(M)\}\subset A_{1}$ and the rows indexed by $B=\{\beta(1),\ldots,\beta(M)\}$. From Theorem 1.7, $g$ is a weighted homogeneous polynomial of weight $w$ and degree $dM-c$, where
$c$ $=\sum_{i=1}^{M}(\alpha(i)-\beta(i))\cdot w$
$=\sum_{i=1}^{M}\alpha(i)\cdot w-\sum_{i=1}^{M}\beta(i)\cdot w.$
We take $C=\mathfrak{A}\setminus C_{n}\subset B_{1}$. We note that $\mathfrak{A}\cap C_{n}\not=\emptyset$ since $\operatorname{Card}(B_{1})=M-1$. Hence $M-(n+1)\leq\operatorname{Card}(C)\leq M-1$. Furthermore, denote $I_{j}=\{(0,1),(0,2),\ldots,(0,j)\}$ for $1\leq j\leq n$, and $C_{i}=\{\beta\in\mathbb{N}^{2}||\beta|=i\}$, for $0\leq i\leq n$.
Proposition 1.14.
Consider the assumptions of Remark 1.13. If $\operatorname{Card}(C)=M-n$, then the following cases hold.
  1. (1)
    If $\mathfrak{A}=(B_{1}\cup C_{n})\setminus I_{n}$, then $c=w_{1}M$.
  2. (2)
    If $\mathfrak{A}\not=(B_{1}\cup C_{n})\setminus I_{n}$, then $|w|M\geq c+|w|$.
Proof.
In order to show (1), we first notice
$\sum_{i=1}^{M}\beta(i)$ $=\sum_{i=0}^{n-1}\left(\sum_{\beta\in C_{i}}\beta\right)$
$=\sum_{i=0}^{n-1}\left(\frac{i(i+1)}{2},\frac{i(i+1)}{2}\right)$
$=\sum_{i=1}^{n}\left(\frac{(i-1)i}{2},\frac{(i-1)i}{2}\right).$
Besides, we note that
$\sum_{i=1}^{M}\alpha(i)$ =$\sum_{i=1}^{n}\left(\sum_{\begin{subarray}{c}\alpha\in C_{i}\\ \alpha\not=(0,i)\end{subarray}}\alpha\right)$
$=\sum_{i=1}^{n}\left(\frac{i(i+1)}{2},\frac{i(i-1)}{2}\right).$
Therefore, $c=\sum_{i=1}^{M}(\alpha(i)-\beta(i))\cdot w=\sum_{i=1}^{n}(i,0)\cdot w=w_{1}M$.
For (2) we consider two cases:
  • $C\not=B_{1}\setminus I_{n-1}$. In this case, there exists $\theta\in B_{1}\setminus C$ such that $\theta_{1}\geq 1$. Set $\mathcal{A}=C$ and so $\operatorname{Card}(B_{1}\setminus\mathcal{A})\geq 2$. Then $|w|M\geq c+|w|$ by Lemma 1.12 (1).
  • $C=B_{1}\setminus I_{n-1}$. In this case $\theta=(1,0)\in C$, and $C_{n}\setminus\mathfrak{A}$ has a unique element $\tau\not=(0,n)$. Then, $\theta\cdot w\leq\tau\cdot w$. Set $\mathcal{A}=C\setminus\{\theta\}$ and so $\operatorname{Card}(B_{1}\setminus\mathcal{A})\geq 2$. Then $|w|M\geq c+|w|$ by Lemma 1.12 (2).
Proposition 1.15.
Consider the assumptions of Remark 1.13. If $\operatorname{Card}(C)\not=M-n$, then $|w|M\geq c+|w|$.
Proof.
We proceed by analyzing all possible cases we can have. Assume $B_{1}=B\setminus\{\beta(1)\}$.
  1. (1)
    $\text{Card}(C)=M-1$. In this case we have $C=B_{1}$. We have that $\theta=(1,0),\theta^{\prime}=(0,1)\in C$. Since $\operatorname{Card}(C_{n}\setminus\mathfrak{A})=n\geq 3$, there exist two vectors $\tau,\tau^{\prime}\in C_{n}\setminus\mathfrak{A}$, $\tau\neq\tau^{\prime}$ such that $\tau_{1},\tau^{\prime}_{2}\geq 1$, so that $\theta\cdot w\leq\tau\cdot w$ and $\theta^{\prime}\cdot w\leq\tau^{\prime}\cdot w$. In this case we take $\mathcal{A}=C\setminus\{\theta,\theta^{\prime}\}$ and the result follows by Lemma 1.12(3).
  2. (2)
    $\text{Card}(C)=M-2$. In this case $C=\mathfrak{A}\setminus C_{n}=B_{1}\setminus\{\beta(j)\}$ for some $j\geq 2$. We have two subcases:
    1. (a)
      $\theta=(1,0)\in C$. In this case, since $\operatorname{Card}(C_{n}\setminus\mathfrak{A})=n-1\geq 2$, there exists $\tau\in C_{n}\setminus\mathfrak{A}$, such that $\tau_{1}\geq 1$ and $\theta\cdot w\leq\tau\cdot w$. We can take $\mathcal{A}=C\setminus\{\theta\}$ and so $\operatorname{Card}(B_{1}\setminus\mathcal{A})\geq 2$. The result follows by Lemma 1.12(2).
    2. (b)
      $(1,0)\not\in C$. We have that $\{(2,0),(1,1)\}\subset C$. As $\operatorname{Card}(C_{n}\setminus\mathfrak{A})=n-1\geq 2$, there exists $\tau\in C_{n}\setminus\mathfrak{A}$ such that $\tau_{1},\tau_{2}\geq 1$ or $\tau=(n,0)$. If $\tau_{1},\tau_{2}\geq 1$ let $\theta=(1,1)$. If $\tau=(n,0)$ let $\theta=(2,0)$. In any of the cases $\theta\in C$ and $\theta\cdot w\leq\tau\cdot w$. We can take $\mathcal{A}=C\setminus\{\theta\}$ and so $\operatorname{Card}(B_{1}\setminus\mathcal{A})\geq 2$. The result follows by Lemma 1.12(2).
  3. (3)
    $\text{Card}(C)=M-j$ for $3\leq j\leq n-1$. We have two subcases:
    1. (a)
      $C=B_{1}\setminus\{(0,b_{1}),...,(0,b_{j-1})\}$. Then $\theta=(1,0)\in C$ and $\operatorname{Card}(C_{n}\setminus\mathfrak{A})=n+1-j\geq 2$. Then there exists $\tau\in C_{n}\setminus\mathfrak{A}$ such that $\theta\cdot w\leq\tau\cdot w$. We can take $\mathcal{A}=C\setminus\{\theta\}$ and so $\operatorname{Card}(B_{1}\setminus\mathcal{A})\geq 2$. The result follows by Lemma 1.12(2).
    2. (b)
      $C=B_{1}\setminus\{(a_{1},b_{1}),...,(a_{j-1},b_{j-1})\}$ with at least one subindex $i$, $1\leq i\leq j-1$ such that $\theta=(a_{i},b_{i})$ satisfies that $a_{i}\geq 1$. In this case, $\theta\in B_{1}\setminus C$. We can take $\mathcal{A}=C$ and so $\operatorname{Card}(B_{1}\setminus\mathcal{A})\geq 2$. The result follows by Lemma 1.12(1).
  4. (4)
    $\operatorname{Card}(C)=M-(n+1)$. In this situation there exists $\theta\in B_{1}\setminus C$ with $\theta_{1}\geq 1$. In this case we can take $\mathcal{A}=C$ and so $\operatorname{Card}(B_{1}\setminus\mathcal{A})\geq 2$. The result follows by Lemma 1.12(1).
Corollary 1.16.
Let $f\in\mathbb{C}[x_{1},x_{2}]$ be a weighted homogeneous polynomial of weight $w\in\mathbb{N}^{2}$ and degree $d$. Suppose that $d\geq 2w_{1}\geq 2w_{2}>0$. Let $n\geq 3$ and $g$ be a maximal minor of $\operatorname{Jac}_{n}(f)$. Then $g$ is a weighted homogeneous polynomial of degree greater or equal than $d$.
Proof.
By Theorem 1.7, $g$ is a weighted homogeneous polynomial of degree $dM-c$. Propositions 1.14 and 1.15 imply $d(M-1)\geq 2w_{1}(M-1)\geq c.$
2. Derivations of the higher Nash blowup local algebra
In this section we study the Lie algebra of derivations of the higher Nash blowup local algebra of a hypersurface defining an isolated singularity.
Definition 2.1.
[HMYZ23, Definition 1.3]. Let $f\in\mathbb{C}\{x_{1},\ldots,x_{s}\}$. Denote
$\mathcal{T}_{n}(f)=\mathbb{C}\{x_{1},\ldots,x_{s}\}/\langle f,\mathcal{J}_{n}(f)\rangle.$
$\mathcal{T}_{n}(f)$ is called the higher Nash blowup local algebra of $f$. For $n=1$ this is also known as the Tjurina algebra of $f$. Moreover, denote as $\mathcal{L}_{n}(f)=\operatorname{Der}(\mathcal{T}_{n}(f))$, i.e., the Lie algebra of derivations of $\mathcal{T}_{n}(f)$.
Remark 2.2.
The higher Nash blowup local algebra is denoted as $\mathcal{M}_{n}$ in [HMYZ23]. In this paper we change the notation in order to be consistent with the usual notation for the Tjurina algebra. In addition, it is also customary to use $\mathcal{M}_{1}$ (or, more precisely, $\mathcal{M}$) for the Milnor algebra.
Let $f$ be a weighted homogeneous polynomial defining an isolated hypersurface singularity. Theorem 1.7 implies that $\mathbb{C}[x_{1},\ldots,x_{s}]/\langle f,\mathcal{J}_{n}(f)\rangle$ is graded. There is an induced grading on $\operatorname{Der}(\mathbb{C}[x_{1},\ldots,x_{s}]/\langle f,\mathcal{J}_{n}(f)\rangle)$, as we explain next.
Remark 2.3.
Suppose that the hypersurface defined by $f$ has an isolated singularity. Then $\mathbb{C}[x_{1},\ldots,x_{s}]/\langle f,\mathcal{J}_{n}(f)\rangle$ is Artinian. This follows from the fact that the zero locus of $\langle f,\mathcal{J}_{n}(f)\rangle$ coincides with the singular locus of the hypersurface [Dua17, Corollary 2.2].
Lemma 2.4 ([XY96, Lemma 2.1]).
Let $A=\bigoplus_{i=0}^{t}A_{i}$ be a graded commutative Artinian local algebra. Let $L(A)$ be the derivation algebra of $A$. Then $L(A)$ can be graded as follows: $L(A)=\bigoplus_{k=-t}^{t}L_{k}$, where $L_{k}=\{D\in L(A)|D(A_{i})\subset A_{i+k}\mbox{ for all }i\}$.
N. Hussain, G. Ma, S. S.-T. Yau, and H. Zuo proposed the following conjecture regarding the non-existence of negative weight derivations on $\mathcal{L}_{n}(f)$.
Conjecture 2.5 ([HMYZ23, Conjecture 1.7]).
Let $f\in\mathbb{C}[x_{1},\ldots,x_{s}]$ be a weighted homogeneous polynomial of weight $w\in\mathbb{N}^{s}$ and degree $d$ that defines an isolated hypersurface singularity. Suppose that $d\geq 2w_{1}\geq\cdots\geq 2w_{s}>0$. Then there is no non-zero negative weight derivation on $\mathcal{T}_{n}(f)$, i.e., $\mathcal{L}_{n}(f)$ is non-negatively graded.
The same authors verified Conjecture 2.5 in the case $s=n=2$ [HMYZ23, Theorem B]. The following theorem gives a positive answer to Conjecture 2.5 for $s=2$ and $n\geq 3$. Together, these results settle the conjecture for two variables.
Theorem 2.6.
Let $f\in\mathbb{C}[x_{1},x_{2}]$ be a weighted homogeneous polynomial of weight $(w_{1},w_{2})\in\mathbb{N}^{2}$ and degree $d$ that defines an isolated hypersurface singularity. Suppose that $d\geq 2w_{1}\geq 2w_{2}>0$. Let $n\geq 3$. Then $\operatorname{Der}(\mathbb{C}[x_{1},x_{2}]/\langle f,\mathcal{J}_{n}(f)\rangle)$ is non-negatively graded.
Proof.
[…]
By the assumptions on the weight and degree of $f$, and the fact that $f$ defines an isolated singularity, we have that $f$ must be one of the following cases [CXY95, Lemma 2.1]:
  • (1)
    $f=x_{1}^{m}+a_{1}(x_{2})x_{1}^{m-1}+\cdots+a_{m}(x_{2})$.
  • (2)
    $f=x_{1}^{m}x_{2}+a_{1}(x_{2})x_{1}^{m-1}+\cdots+a_{m}(x_{2})$.
By Corollary 1.16, $f$ is a generator of smallest degree among the generators of $\langle f,\mathcal{J}_{n}(f)\rangle$. Since $D(f)\in\langle f,\mathcal{J}_{n}(f)\rangle$ and $D$ is of negative degree, it follows that $D(f)=0$, i.e., $cx_{2}^{b}\partial_{x_{1}}(f)=0$. Using (1) and (2) we obtain $\partial_{x_{1}}(f)\neq 0$. We conclude that $c=0$ and so $D=0$. ∎
Remark 2.7.
Consider the notation of Theorem 2.6. If $f\in\langle x_{1},x_{2}\rangle^{3}$, then we can assume $d\geq 2w_{1}\geq 2w_{2}>0$ without loss of generality [YZ16b, Theorem 2.1], [Sai71].
Acknowledgements
The first author was supported by UNAM Posdoctoral Program (POSDOC). The second author was supported with project IN100723, “Curvas, Sistemas lineales en superficies proyectivas y fibrados vectoriales? from DGAPA, UNAM. The third author is supported by CONAHCYT project CF-2023-G-33.